OpenStudy (anonymous):

Test the polar equation for symmetry with respect to the polar axis, the pole, and the line θ = π/2 r = 4 cos θ csc θ

OpenStudy (anonymous):

to test with respect to the polar axis, set theta to negative theta

OpenStudy (anonymous):

\[4\cos (-\theta)\csc (-\theta)\]

OpenStudy (anonymous):

-4\[-4\cos (\theta)-\csc (\theta)\]